# Engineering impurity Bell states through coupling with a quantum bath

### v1

@@ IV Improving impurity-impurity correlations through environment engineering
As we have shown, the coupling to the C component plays a crucial in spatially separating the impurities from the bath and being able to create the impurity Bell states. On the other hand, the impurity-bath correlations that are created by this coupling also negatively affects the purity of the entangled impurity states and consequently cannot be completely separated from the bath. However, the properties of the bath are also tunable, for instance, by choosing a different atomic species the mass ratio $m_{\text{C}}/m$ can be varied, while there is also some degree of control over the intraspecies interactions $g_{\rm{C}}$. Let us first focus on adjusting the mass ratio while keeping $g_{\rm{C}}=0$ fixed. The von Neumann entropy, concurrence, mutual information and binding energy are shown in Fig. 4 as a function of the mass ratio for fixed impurity-impurity interactions $g_{\mathrm{AB}}=-3,0,3$.
It is immediately apparent from looking at the decay of the von Neumann entropy and tripartite mutual information that increasing the mass of the bath particles can significantly reduce the degree of the impurity-bath correlations. This can be well described by an algebraic decay of the form $S^{\text{vN}}_{\text{AB}}\sim\left(m_{\text{C}}/m\right)^{-\alpha}$ and we show the fitted decay rates in Table 1. The same effect, albeit with a smaller decay rate, can be achieved by increasing the number of particles $N_{\text{C}}$, while keeping $m_{\text{C}}/m=1$ (see inset of Fig. 4(b)), with the fits to $S^{\text{vN}}_{\text{AB}}\sim N_{\text{C}}^{-\beta}$ shown again in Table 1. It is evident that the exponent $\alpha$ is almost unaffected by the interaction strength between the two impurities, while the decay rate $\beta$ can be increased in the regime of repulsive impurity-impurity interactions. We also note that there is a minimum number of C particles that are needed to spatially separate the impurities, for instance $N_{\text{C}}>3$ for $g_{\text{AB}}=3$, as for smaller bath sizes the impurities are bunched in the center of the trap and hence the concurrence is zero and the von Neumann entropy takes significantly larger values.
@@
Finally, in Fig. 5 we examine the effect of the bath intraspecies coupling strength, $g_{\rm{C}}$, on the concurrence and von Neumann entropy for fixed $m_{\rm{C}}=m$. Increasing the repulsive interactions between the C component atoms broadens the density distribution of the bath, and leads to an increase in the impurity-bath correlations due to larger wave-function overlap in the finite trap environment. This also adversely affects the concurrence for all the three cases we consider, with the sudden decrease for the case $g_{\text{AB}}=3$ again a signal of this state transitioning from being anti-bunched to bunched. On the other hand, for attractive interactions $g_{\rm{C}}<0$ the concurrence is increased and the von Neumann entropy reduced, which indicates better conditions for the formation of the $|\Phi^{+}\rangle$ Bell state. Since the attractively interacting environment plays a role similar to the one of increasing mass, albeit to a lesser extent (see green dash-dotted lines in Fig. 3 and $\rho_{\text{LR}}$ in Fig. 6(b)), a combination of both effects can be used to enhance the impurity-impurity correlations. To highlight this we show in Fig. 3 how bath interactions $g_{\text{C}}=-5$ in combination with a mass ratio of $m_{\text{C}}=15m$ (yellow dashed lines), can be used to significantly improve the impurity-impurity correlations over the whole range of $g_{\text{AB}}$. Specifically, for $g_{\text{AB}}=-5$ the concurrence reaches $\mathcal{C}\approx 0.964$ while the von Neumann entropy is $S^{\rm{vN}}_{\rm{AB}}\approx 0.125$. This characterises a state that is close to the $|\Phi^{+}\rangle$ Bell state and we show $\rho_{\text{LR}}$ for this case in Fig. 6(c). We also note that the TMI is negative for $g_{\text{AB}}<0$, indicating that the composite system AB has more information compared to the individual systems A and B [52] due to it forming a tightly bound state.
@@ V Conclusions
To summarize, we have proposed a scheme for robust preparation of strongly-correlated states with current ultra-cold atomic setups, where particle numbers and interaction strengths are experimentally controllable, and where the spatial correlations we describe can be easily measured. We numerically demonstrate that strongly correlated states close to Bell states can be formed as ground states in systems of two distinguishable impurities immersed in a bosonic reservoir with a well-defined particle number. Our analysis shows that the Bell state $|\Psi^{+}\rangle$ can be readily generated with two repulsively interacting impurities, while the Bell state $|\Phi^{+}\rangle$ is hard to achieve due to large correlations created between the bath and impurities which reduce the states coherence. With the aim of reducing the impact of impurity-bath correlations such that the entanglement between the two impurities can be further enhanced, we have investigated the properties of the bath including its intra-species coupling strength, and the mass of its particles. We have demonstrated that in both situations the formation of the Bell states is substantially enhanced compared to the mass-balanced non-interacting bath. Importantly, we have shown in this work that the bosonic bath located in the center of the harmonic trap forms a matter-barrier to separate the two impurities resulting in the emergence of their non-classical properties. We emphasize that this effect is a purely quantum phenomenon proven by the finite von Neumann entropy between the bosonic bath and impurities and that it is significantly different from classical potentials. We anticipate that our results not only pave an efficient and experimentally feasible way to create and fully control strongly correlated atomic states, but also will stimulate further research on the non-classical properties of multi-component quantum systems.

### v2

@@ IV Improving impurity-impurity correlations through environment engineering
As we have shown, the coupling to the C component plays a crucial role in spatially separating the impurities from the bath and being able to create the impurity Bell states. On the other hand, the impurity-bath correlations that are created by this coupling also negatively affect the purity of the entangled impurity states and consequently cannot be completely separated from the bath. However, the properties of the bath are also tunable; for instance, by choosing a different atomic species the mass ratio $m_{\text{C}}/m$ can be varied, while there is also some degree of control over the intraspecies interactions $g_{\rm{C}}$. Let us first focus on adjusting the mass ratio while keeping $g_{\rm{C}}=0$ fixed. The von Neumann entropy, concurrence, mutual information and binding energy are shown in Fig. 4 as a function of the mass ratio for fixed impurity-impurity interactions $g_{\mathrm{AB}}=-3,0,3$.
It is immediately apparent from looking at the decay of the von Neumann entropy and tripartite mutual information that increasing the mass of the bath particles can significantly reduce the degree of the impurity-bath correlations. This can be well described by an algebraic decay of the form $S^{\text{vN}}_{\text{AB}}\sim\left(m_{\text{C}}/m\right)^{-\alpha}$ and we show the fitted decay rates in Table 1. It is worth mentioning that the decrease of the bath-mediated attractive interactions due to the larger mass of the bath particles compared to the impurities is in good agreement with the prediction in Ref. [22]. The same effect, albeit with a smaller decay rate, can be achieved by increasing the number of particles $N_{\text{C}}$, while keeping $m_{\text{C}}/m=1$ (see inset of Fig. 4(b)), with the fits to $S^{\text{vN}}_{\text{AB}}\sim N_{\text{C}}^{-\beta}$ shown again in Table 1. It is evident that the exponent $\alpha$ is almost unaffected by the interaction strength between the two impurities, while the decay rate $\beta$ can be increased in the regime of repulsive impurity-impurity interactions. We also note that there is a minimum number of C particles that are needed to spatially separate the impurities, for instance $N_{\text{C}}>3$ for $g_{\text{AB}}=3$, as for smaller bath sizes the impurities are bunched in the center of the trap and hence the concurrence is roughly zero and the von Neumann entropy takes significantly larger values as shown in the insets in panel (a) and (b) in Fig. 4.
@@
Figure 7: The Bell states $|\Psi^{-}\rangle$ and $|\Phi^{-}\rangle$ are the first excited state of the corresponding system whose ground state demonstrates the states $|\Psi^{+}\rangle$ and $|\Phi^{+}\rangle$. The four columns are the OBDM $\rho^{(1)}_{\rm{A}}(x,x^{\prime})$, the OBDM $\rho^{(1)}_{\rm{C}}(x,x^{\prime})$, the diagonal TBDM $\rho^{(2)}_{\rm{AB}}(x_{\rm{A}},x_{\rm{B}})=\rho^{(2)}_{\rm{AB}}(x_{\rm{A}},x_{\rm{B}},x_{\rm{A}},x_{\rm{B}})$, and the reduced density matrix $\rho_{\rm{LR}}$, respectively. The different rows correspond to different cases, $g_{\rm{AB}}=-5$, $m_{\rm{C}}/m=15$ (upper row), and $g_{\rm{AB}}=20$, $m_{\rm{C}}/m=1$ (lower row). In all panels, the bosonic bath has $N_{\rm{C}}=10$ non-interacting ($g_{\rm{C}}=0$) bosons, while the impurity-bath coupling strengths are kept fixed $g_{\rm{AC}}=g_{\rm{BC}}=5$. Note that all density matrices are normalized to unity.
Finally, in Fig. 5 we examine the effect of the bath intraspecies coupling strength, $g_{\rm{C}}$, on the concurrence and von Neumann entropy for fixed $m_{\rm{C}}=m$. Increasing the repulsive interactions between the C component atoms broadens the density distribution of the bath, and leads to an increase in the impurity-bath correlations due to larger wave-function overlap in the finite trap environment. This also adversely affects the concurrence for all the three cases we consider, with the sudden decrease for the case $g_{\text{AB}}=3$ again a signal of this state transitioning from being anti-bunched to bunched. On the other hand, for attractive interactions $g_{\rm{C}}<0$ the concurrence is increased and the von Neumann entropy reduced, which indicates better conditions for the formation of the $|\Phi^{+}\rangle$ Bell state. Since the attractively interacting environment plays a role similar to the one of increasing mass, albeit to a lesser extent (see green dash-dotted lines in Fig. 3 and $\rho_{\text{LR}}$ in Fig. 6(b)), a combination of both effects can be used to enhance the impurity-impurity correlations. To highlight this we show in Fig. 3 how bath interactions $g_{\text{C}}=-5$ in combination with a mass ratio of $m_{\text{C}}=15m$ (yellow dashed lines), can be used to significantly improve the impurity-impurity correlations over the whole range of $g_{\text{AB}}$. Specifically, for $g_{\text{AB}}=-5$ the concurrence reaches $\mathcal{C}\approx 0.964$ while the von Neumann entropy is $S^{\rm{vN}}_{\rm{AB}}\approx 0.125$. This characterizes a state that is close to the $|\Phi^{+}\rangle$ Bell state and we show $\rho_{\text{LR}}$ for this case in Fig. 6(c). We also note that the TMI is negative for $g_{\text{AB}}<0$, indicating that the composite system AB has more information compared to the individual systems A and B [55] due to it forming a tightly bound state.
It is also worth noting that the spatially anti-symmetric Bell states, $|\Psi^{-}\rangle=\left(|\rm{LR}\rangle-|\rm{RL}\rangle\right)/\sqrt{2}$ and $|\Phi^{-}\rangle=\left(|\rm{LL}\rangle-|\rm{RR}\rangle\right)/\sqrt{2}$, naturally appear as the first excited states of the system. Although the OBDM and the diagonal TBDM of the first excited state are quite similar to those of the ground state due to the fact that they are nearly degenerate, we show in Fig. 7 that a clear difference appears in the reduced density matrix in the $\{|\text{L}\rangle,|\text{R}\rangle\}$ basis $\rho_{\text{LR}}$ as the coherence terms are necessarily negative. Furthermore, we note that like the spatially symmetric Bell states $|\Psi^{+}\rangle$ and $|\Phi^{+}\rangle$, we can also improve the fidelity of $|\Psi^{-}\rangle$ and $|\Phi^{-}\rangle$ by modifying the bath parameters such as its mass, giving us qualitatively similar results as discussed previously.
@@ V Conclusions
To summarize, we have proposed a scheme for robust preparation of strongly-correlated states with current ultra-cold atomic setups, where particle numbers and interaction strengths are experimentally controllable, and where the spatial correlations we describe can be easily measured. We numerically demonstrate that strongly correlated states close to Bell states can be formed as ground states in systems of two distinguishable impurities immersed in a bosonic reservoir with a well-defined particle number. Our analysis shows that the Bell state $|\Psi^{+}\rangle$ can be readily generated with two repulsively interacting impurities, while the Bell state $|\Phi^{+}\rangle$ is hard to achieve due to large correlations created between the bath and impurities which reduce the states coherence. With the aim of reducing the impact of impurity-bath correlations such that the entanglement between the two impurities can be further enhanced, we have investigated the properties of the bath including its intra-species coupling strength, and the mass of its particles. We have demonstrated that in both situations the formation of the Bell states is substantially enhanced compared to the mass-balanced non-interacting bath. Importantly, we have shown in this paper that the bosonic bath located in the center of the harmonic trap forms a matter-barrier to separate the two impurities resulting in the emergence of their non-classical properties. We emphasize that this effect is a purely quantum phenomenon proven by the finite von Neumann entropy between the bosonic bath and impurities and that it is significantly different from classical potentials. We anticipate that our results not only pave an efficient and experimentally feasible way to create and fully control strongly correlated atomic states, but also will stimulate further research on the non-classical properties of multi-component quantum systems.
@@ VI Acknowledgments
Appendix A Numerical Method
In the following, we briefly summary the ab initio method used in this work, the improved Exact Diagonalization method [44]. For numerical purposes, it is naturally convenient to diagonalize the many-body Hamiltonian in the formalism of second quantization. In doing so, the field operators are introduced as
$\hat{\Psi}_{\sigma}(x)=\sum\limits_{k}\phi_{\sigma,k}(x)\hat{a}_{\sigma,k},$ (10)
$\hat{\Psi}_{\sigma}^{\dagger}(x)=\sum\limits_{k}\phi^{*}_{\sigma,k}(x)\hat{a}^{\dagger}_{\sigma,k}.$ (11)
The field operator $\hat{\Psi}_{\sigma}(x)$ ($\hat{\Psi}^{\dagger}_{\sigma}(x)$) annihilates (creates) a $\sigma$-species boson being in the single-particle state $\phi_{\sigma,k}(x)$ at position $x$. Here $\hat{a}_{\sigma,k}$ ($\hat{a}^{\dagger}_{\sigma,k}$) corresponds to the annihilation (creation) operator. For bosonic systems, the creation and annihilation operators must obey the commutation relations
$\left[\hat{a}_{\sigma,k},\hat{a}_{\delta,\ell}^{\dagger}\right]$ $=\delta_{\sigma\delta}\delta_{k\ell},$ (12)
$\left[\hat{a}_{\sigma,k}^{\dagger},\hat{a}_{\delta,\ell}^{\dagger}\right]$ $=\left[\hat{a}_{\sigma,k},\hat{a}_{\delta,\ell}\right]=0.$ (13)
The resulting Hamiltonian is then given as
$\hat{H}=$ $\sum_{\sigma\in\{A,B,C\}}\sum_{k,\ell}h^{\sigma}_{k\ell}\hat{a}^{\dagger}_{\sigma,k}\hat{a}_{\sigma,\ell}$
$+\dfrac{1}{2}\sum_{k\ell mn}W^{\text{C}}_{k\ell mn}\hat{a}^{\dagger}_{\text{C},k}\hat{a}^{\dagger}_{\text{C},\ell}\hat{a}_{\text{C},m}\hat{a}_{\text{C},n}$
$+\sum_{\sigma\neq\delta\in\{A,B,C\}}\sum_{k\ell mn}W^{\sigma\delta}_{k\ell mn}\hat{a}^{\dagger}_{\sigma,k}\hat{a}^{\dagger}_{\delta,\ell}\hat{a}_{\sigma,m}\hat{a}_{\delta,n}.$ (14)
Here $h^{\sigma}_{k\ell}$, $W^{\text{C}}_{k\ell mn}$, and $W^{\sigma\delta}_{k\ell mn}$ are one- and two-body matrix elements expressed in the basis of the single-particle Hamiltonian. As the system we consider is confined in a 1D harmonic trap, it is straightforward to employ the harmonic oscillator eigenfunctions as the single-particle functions $\phi_{\sigma,k}(x)$. This choice makes the uses of analytical results to obtain the matrix elements $h^{\sigma}_{k\ell}=\left(k+0.5\right)\delta_{k\ell}$, $W^{\text{C}}_{k\ell mn}$ and $W^{\sigma\delta}_{k\ell mn}$, thus the convergence is accelerated. Further details of this approach can be found in Refs. [58, 59, 60, 44, 61].
The ansatz wavefunction is factorized as a linear combination of a set of orthonormal Fock states
$|\Psi\rangle=\sum_{j_{\rm{A}}=1}^{D_{\text{A}}}\sum_{j_{\rm{B}}=1}^{D_{\text{B}}}\sum_{j_{\rm{C}}=1}^{D_{\text{C}}}c_{j_{\rm{A}},j_{\rm{B}},j_{\rm{C}}}|\Phi^{\text{A}}_{j_{\rm{A}}}\rangle|\Phi^{\text{B}}_{j_{\rm{B}}}\rangle|\Phi^{\text{C}}_{j_{\rm{C}}}\rangle.$ (15)
Here, $c_{j_{\rm{A}},j_{\rm{B}},j_{\rm{C}}}$ denote the expansion coefficients and $|\Phi^{\sigma}_{j_{\sigma}}\rangle=|n^{\sigma}_{1},n^{\sigma}_{2}\dots n^{\sigma}_{k}\dots\rangle$ is a possible $\sigma$-species permanent (also known as configuration) in the total of $D_{\sigma}$ permanents used to expand the ansatz. Note that in each permanent, the occupation number $n^{\sigma}_{k}$ in the single-particle state $\phi_{\sigma,k}(x)$ can be arbitrary integers between 0 and $N_{\sigma}$ and must obey the constraint $\sum\limits_{k}n_{k}^{\sigma}=N_{\sigma}$. In practice, the permanents for the ansatz are selected such that their energies in the non-interacting many-body Hamiltonian are less than a certain optimal value $E_{max}$. This truncation approach is known as the energy-cutoff scheme proposed in Refs. [62, 63] and hence the accuracy of the numerical results is controlled by increasing $E_{max}$. The key ingredient of the improved ED scheme which significantly reduces the number of required permanents is the selection of dominant configurations in terms of the spatial symmetry of the desired many-body state [44]. It is known that if the trapping potential of a quantum particle has a spatial symmetry, namely
$V(x)=V(-x),$ (16)
then the single-particle eigenfunctions $\phi_{k}(x)$ possess a well-defined spatial symmetry. Mathematically, the spatially symmetric $\phi_{k=2n}(x)$ are even functions, while the spatially asymmetric $\phi_{k=2n+1}(x)$ are odd functions. Because the permanents are the symmetrized Hartree product of the single-particle functions $\phi_{k}(x)$, they must satisfy this spatial symmetry. This leads to the fact that the desired wave functions are solely expanded by either even- or odd-parity permanents. Thus, it is more practical to use only the permanents that have the same parity as the desired many-body wavefunction in the expansion of the ansatz. Since the bosonic ground-state wavefunction is spatially symmetric, we therefore only use even-parity permanents for the ansatz in this work. Meanwhile, for obtaining the first excited state, the ansatz is expanded by odd-parity permanents.
In general, minimizing the expectation value of the Hamiltonian (A) with respect to the ansatz (15) leads to standard Hermitian eigenvalue problem which can be written in the form of
$\mathbf{H}|C_{m}\rangle=E_{m}|C_{m}\rangle,$ (17)
with $\mathbf{H}$ being the matrix representation of the many-body Hamiltonian (A). The pair $\{E_{m},|C_{m}\rangle\}$ is the $m$-th many-body eigenvalue and eigenvector, respectively.